### v1

# Probing non-Markovian qubit noise and modeling Post Markovian Master Equation

Abstract
Understanding the noise characteristics of quantum processors is crucial when achieving fault-tolerant quantum computing. However, typical qubit designs are often studied under the Markovian approximation, which does not fully capture realistic dynamics. Factors such as qubit-qubit coupling and extended bath correlation times can introduce significant non-Markovian effects into the noise processes. In this study, we employ the Post-Markovian Master Equation (PMME) formalism to characterize memory effects in the noise dynamics. We further experimentally validate the PMME framework using superconducting qubits on an IBM Quantum device, demonstrating clear non-Markovian behavior during circuit execution. Additionally, we quantify the crosstalk effect using an information-theoretic approach and reveal that crosstalk can dominate the observed non-Markovian effects in current quantum hardware.
I Introduction
The dynamics of contemporary quantum processors is shaped by environmental fluctuations that are often temporally correlated. While the Markov approximation leads to a time-homogeneous Lindblad semigroup and offers a tractable model for control and error analysis, experiments across solid-state platforms—and especially in superconducting qubits—display correlated noise (e.g., $1/f$ spectra), long-lived crosstalk, nonexponential relaxation, and coherence revivals that invalidate a purely Markovian description [1, 2, 3, 4]. A compact, operational framework that both diagnoses and models memory is therefore valuable for device characterization, calibration, and scheduling.
Non-Markovianity and operational diagnostics.
Several complementary notions formalize quantum non-Markovianity. In the information-backflow picture, revivals of the trace distance between two system states certify non-Markovian dynamics [5]. In the CP-divisibility picture, non-Markovianity manifests as the loss of complete positivity of intermediate maps $V_{t,s}=\Phi_{t}\Phi_{s}^{-1}$, $t>s$ [6]. Both criteria are experimentally meaningful once one reconstructs dynamical maps $\Phi_{t}$ via quantum process tomography (QPT) [7, 8].
From projection-operator formalisms to PMME.
Projection-operator methods provide two exact yet generally intractable descriptions: the time-nonlocal Nakajima–Zwanzig (NZ) memory-kernel equation and the time-local time-convolutionless (TCL) expansion [9, 10, 11]. To interpolate between the analytically friendly Lindblad limit and fully non-Markovian kernels, Shabani and Lidar introduced the post-Markovian master equation (PMME), which preserves complete positivity under mild conditions while incorporating bath memory through a scalar kernel [12]. In its widely used form,
$\frac{d}{dt}\rho(t)=\mathcal{L}\!\int_{0}^{t}\!k(\tau)\,e^{\mathcal{L}\tau}\,\rho(t-\tau)\,d\tau,$ (1)
where $\mathcal{L}$ is a (Markovian) Lindblad generator and $k(\tau)$ is a phenomenological memory kernel. Equation (70) reduces to Lindblad in the $\delta$-kernel limit, $k(\tau)=\gamma\,\delta(\tau)$, and corresponds to a translation-invariant NZ kernel of the form $\mathsf{K}(\tau)=k(\tau)\,\mathcal{L}e^{\mathcal{L}\tau}$ [12]. Solving (70) in the damping basis of $\mathcal{L}$ yields closed expressions in Laplace space, enabling data-driven kernel reconstruction.
Superconducting qubits as a testbed.
In transmon architectures, slow material noise and parasitic ZZZZ couplings generate device-level memory that impacts idle evolution and concurrent gates. Experimental studies report coherence revivals and extra spectral features diagnostic of non-Markovianity; targeted control (tunable couplers, dynamical decoupling) suppresses ZZZZ crosstalk but does not eliminate all memory [13, 14, 15, 4]. A channel-level summary that (i) detects non-Markovianity directly from tomographic data and (ii) models the same data via a physically interpretable kernel $k(t)$ is therefore practically useful.
This work: tomography $\rightarrow$ witnesses $\rightarrow$ PMME kernel.
We develop and deploy an end-to-end pipeline on IBM-style superconducting hardware. First, we reconstruct time-indexed dynamical maps $\{\Phi_{t}\}$ via QPT. Second, we diagnose non-Markovianity using both operational witnesses: (a) information-backflow via trace-distance revivals, and (b) CP-divisibility violations detected by Choi positivity of $V_{t,s}=\Phi_{t}\Phi_{s}^{-1}$. Third, we fit a best-fit Lindblad envelope $\mathcal{L}$ capturing the long-time decay and then reconstruct the post-Markovian kernel $k(t)$ by working in the damping basis of $\mathcal{L}$ and numerically inverting the Laplace-domain solution implied by (70). Across multiple, distinct initial conditions we observe (i) extended CP-divisibility violations, (ii) trace-distance revivals, and (iii) an oscillatory, approximately state-independent kernel $k(t)$ that captures the short-time departures from the Lindblad envelope. These results provide a compact, interpretable summary of device memory that directly informs mitigation (layout-aware scheduling and decoupling) and model-based simulation.
II Introduction
In the ideal model of quantum computation, gate operations are always assumed to be perfectly unitary. However, in the physical implementation, a qubit is constructed by physical components and is impossible to fully isolate from the rest of the world. The device of a quantum processor is intrinsically open and coupled to the underlying environment. As a result, the dynamics will have the mixing with the noise process originating from the coupling to the environment. Thus, the study of Open Quantum Systems (OQS) has been a central topic in quantum information science. The ability to study and characterize the noise model of a given system is not simply a theoretical interest but a practical necessity, indispensable on the road to fault-tolerant quantum computation [16].
As mentioned by Shabani and Lidar in 2005 [17] An open quantum system coupled with Hilbert Spaces (HS, HB), which represent System environment and Bath environment, and these two space build the system-bath Hamiltonian HSB which could represent as:
$\rho(t)=\mathrm{Tr}_{B}[U(t)\rho_{SB}(0)U^{\dagger}(t)],$ (2)
A widely used approach to modeling noise in quantum systems is the Lindblad master equation, which assumes that the environment is memoryless (Markovian). The Markovian approximation allows for an accurate description of exponential decay processes such as faithfully recovering the excited-state lifetime (T1) and coherence time (T2) [18]. However, in real-world quantum processors, the Markovian assumption often breaks down. Long bath correlation times, crosstalk between qubits, and strong coupling to the environment introduce memory effects leading to non-Markovian noise behavior [19]. Such noise manifests as deviations from simple exponential decay such as oscillatory or non-exponential relaxation patterns that challenge the traditional error correction strategies. Indeed, recent studies have demonstrated that non-Markovian noise can significantly lower error-correction thresholds, implying that the presence of memory effects can drastically impact the performance of error-corrected quantum devices [20]. Therefore, it becomes imperative to investigate these non-Markovian characteristics and develop methods that mitigate their detrimental effects on quantum computation.
Among the various approaches to model non-Markovian dynamics, the Post-Markovian Master Equation (PMME) offers a mathematically manageable and experimentally relevant framework for doing so. Critically, it introduces a memory kernel function $k(\tau)$ that characterizes the degree and timescale of memory effects. The equation reads:
$\dot{\rho}(t)=\mathcal{L}_{0}\rho(t)+\mathcal{L}_{1}\int_{0}^{t}d\tau e^{(\mathcal{L}_{0}+\mathcal{L}_{1})\tau}k(\tau)\rho(t-\tau)$ (3)
Here, $\mathcal{L}_{0}$ represents the Markovian (memory-less) component, while $\mathcal{L}_{1}$ governs the non-Markovian contribution, weighted by the memory kernel. In the limit where $k(\tau)=\delta(\tau)$, the equation reduces to the conventional Lindblad form, indicating purely Markovian dynamics. Our preliminary experiments on IBM quantum hardware achieved an noticeable behavior. Our first experiment tested long-time behavior by setting 10000 50000 idling gates. Since the gate operations for the IBM quantum machines is around 68 ns [21], the total running time was over 0.68 ms. We found that when the observation time is relatively long, the system exhibits Markovian behavior evidenced by exponential coherence decay, but shortening the observation time reveals signatures of non-Markovian noise consistent with findings in recent literature. This indicates that memory effects become more prominent at shorter timescales, suggesting that the noise model’s nature may depend on the temporal resolution of the measurements. Notably, similar results were reported by Tripathi et al. [22] where non-Markovian characteristics were observed in short-time dynamics despite a seemingly Markovian long-time behavior.
Figure 1: Population decay for 50000 ID gates
Figure 2: Bloch vector under 5000 ID gates
The choice to use entangled states is due to their inherent sensitivity to environmental disturbances. Focusing on the quantum coherence through time evolution allows one to detect non-Markovian noise. Indeed, recent study have shown that non-Markovian noise can significantly destabilize entangled states, making them ideal for characterizing complex noise behaviors [23]. For this paper, we utilized a two-qubit entangled state prepared in the ++ configuration on an IBM quantum processor followed by full-state tomography to assess the noise characteristics.
In this paper, we experimentally characterize non-Markovian noise on IBM quantum devices by leveraging the PMME framework. We aim to systematically investigate the role of short-term versus long-term correlations in noise processes, using two-qubit entanglement as a diagnostic tool. By obtaining a deeper understanding of how non-Markovian noise influences coherence times, we seek to develop enhanced error-correction strategies tailored to practical quantum processors.
III Preliminaries
In this section, we review the Post-Markovian Master Equation (PMME) formalism and introduce information-theoretic methods for quantifying both qubit crosstalk and non-Markovian noise effects.
III.1 Non-Markovianity via CP-Divisibility
A rigorous way to characterize non-Markovian dynamics employs the notion of CP-divisibility. A one-parameter quantum map $\Phi_{t}$ is said to be CP-divisible if, for all $t\geq s\geq 0$, there exists a completely positive and trace-preserving (CPTP) intermediate map $\mathcal{V}_{t,s}$ such that
$\Phi_{t}=\mathcal{V}_{t,s}\circ\Phi_{s}.$ (4)
The evolution is Markovian if and only if the map is CP-divisible for all $t\geq s$. Lack of CP-divisibility thus signals non-Markovian behavior.
To test CP-divisibility, one can inspect the Choi matrix
$\chi_{t}=(\mathcal{I}\otimes\Phi_{t})(|\gamma\rangle\langle\gamma|),$ (5)
where $|\gamma\rangle=\sum_{i}|i\rangle\otimes|i\rangle$ is the (unnormalized) maximally entangled state on $\mathcal{H}\otimes\mathcal{H}$. A map $\Phi_{t}$ is CP if and only if $\chi_{t}$ is positive semidefinite ($\chi_{t}\geq 0$).
Directly obtaining the Choi matrix for the intermediate map $\mathcal{V}_{t,s}$ is generally infeasible, since the system and bath cannot be reinitialized at time $s$. Instead, one may reconstruct the dynamical map in the operator-sum (superoperator) representation. By the Choi–Jamiołkowski isomorphism, any quantum map satisfies
$\Phi_{t}(\rho)=\mathrm{Tr}_{1}\bigl[\chi_{t}(\rho^{T}\otimes I)\bigr].$ (6)
If we vectorize the density operator as $\mathrm{vec}(\rho)$, the map can be expressed as
$\mathrm{vec}\bigl(\Phi_{t}(\rho)\bigr)=S_{t}\mathrm{vec}(\rho),$ (7)
where the superoperator matrix $S_{t}$ and the Choi matrix $\chi_{t}$ are related by
$(S_{t})_{ij,k\ell}=(\chi_{t})_{ki,\ell j}.$ (8)
which is an index reshuffling map, the detailed derivation can be found in Appendix A. Therefore, we can write $S_{t}=\mathcal{R}(\chi_{t})$, where $\mathcal{R}$ is the permutation map. Therefore, the Choi matrix for $\mathcal{V}_{t,s}$ can be written as
$\chi_{t,s}=\mathcal{R}^{-1}(S_{t}S_{s}^{-1})$ (9)
By reconstructing $\chi_{t}$ (and thus $S_{t}$) via quantum process tomography at different times, one can infer the complete positivity of the Choi matrix of the intermediate map and thus diagnose CP-divisibility (and hence non-Markovianity) indirectly.
III.2 Non-Markovianity via Information-Backflow
Another way we characterize non-Markovianity is by looking into the information backflow. In the Markovian case, the quantum information flows monotonically to the environment, and this phenomenon can be characterizing by the distinguishability between states. The most common distinguishability measures are trace distance and quantum relative entropy:
$T(\rho,\sigma)$ $\coloneqq\frac{1}{2}\mathrm{Tr}|\rho-\sigma|$ (10)
$D(\rho\|\sigma)$ $\coloneqq\mathrm{Tr}\rho(\log\rho-\log\sigma)$ (11)
and are a metric for how well two states can be distinguished using measurement. An important and fundamental property in quantum information theory is the Data Processing Inequality for trace distance and quantum relative entropy, which states that if a map $\Phi$ is CPTP then
$T(\Phi_{t}(\rho),\Phi_{t}(\sigma))$ $\leq T(\rho,\sigma)$ (12)
$D(\Phi_{t}(\rho),\Phi_{t}(\sigma))$ $\leq D(\rho,\sigma)$ (13)
and the two quantities are contractive under $\Phi_{t}$. We can now define a new measure that is the distinguishability across time of two initial states evolving under a quantum map $\Phi_{t}$. This is given as
$D(t,\rho_{1,2})=\frac{1}{2}\mathrm{Tr}|\Phi_{t}(\rho_{1})-\Phi_{t}(\rho_{2})|.$ (14)
For a CPTP map, $\sigma(t,\rho_{1,2})=\frac{d}{dt}D(t,\rho_{1,2})<0\forall t$. Since the map is contractile and losing information at all times and for all initial states, it is Markovian. If for any t and for any pair of $\rho_{1,2}$ the aforementioned inequality does not hold, information is flowing back into the system from the environment. This means that there exists some memory of previous states generating non-Markovian evolution. Finding $D(t,\rho_{1,2})$ involves simply measuring $\rho$ through time and calculating $\sigma$. If $\sigma$ ever has a value greater than 0, the evolution is non-Markovian.
III.3 Information‑theoretic probe of crosstalk
An intuitive signature of crosstalk is the emergence of correlations between two qubits that were initially uncorrelated. Let the qubits AA and $B$ be prepared in a product state $\rho_{A}\otimes\rho_{B}$. After a joint noise process, the state evolves to a (generically) correlated density matrix $\rho_{AB}(t)$. The total correlation generated can be quantified by the quantum mutual information
$I(A\!\mathrel{\mathop{\ordinarycolon}}\!B)_{\rho(t)}=S\!\bigl(\rho_{A}(t)\bigr)+S\!\bigl(\rho_{B}(t)\bigr)-S\!\bigl(\rho_{AB}(t)\bigr)$ (15)
where $S(\cdot)$ denotes the von Neumann entropy and $\rho_{A(B)}(t)=\operatorname{Tr}_{B(A)}[\rho_{AB}(t)]$. A mathematically equivalent expression that highlights the “distance” from a product state is
$I(A\!\mathrel{\mathop{\ordinarycolon}}\!B)_{\rho(t)}=D\!\bigl(\rho_{AB}(t)\,\bigl\|\,\rho_{A}(t)\otimes\rho_{B}(t)\bigr),$ (16)
with $D(\rho\|\sigma)\!=\!\operatorname{Tr}\!\bigl[\rho\bigl(\log\rho-\log\sigma\bigr)\bigr]$ being the quantum relative entropy. Hence, $I(A\!\mathrel{\mathop{\ordinarycolon}}\!B)_{\rho(t)}=0$ if and only if $\rho_{AB}(t)$ factorizes into $\rho_{A}(t)\otimes\rho_{B}(t)$, i.e. the qubits remain completely uncorrelated. Any positive value therefore witnesses crosstalk‑induced correlations—whether classical, quantum, or both. In practice we tomographically reconstruct $\rho_{AB}(t)$ at discrete idle times and track the growth (or decay) of $I(A\!\mathrel{\mathop{\ordinarycolon}}\!B)_{\rho(t)}$ as a sensitive, information‑theoretic figure of merit for device‑level crosstalk.
III.4 Post Markovian Master Equation
We start from the Gorini–Kossakowski–Sudarshan–Lindblad generator
$\mathcal{L}(\rho)$ $=\frac{i\omega_{z}}{2}\,[\sigma_{z},\rho]+\gamma_{\mathrm{AD}}\!\Bigl(\sigma_{-}\rho\sigma_{+}-\tfrac{1}{2}\{\sigma_{+}\sigma_{-},\rho\}\Bigr)$
$\hskip 56.9055pt+\gamma_{\mathrm{PD}}\!\bigl(\sigma_{z}\rho\sigma_{z}-\rho\bigr),$ (17)
where $\omega_{z}$ is the qubit precession frequency, $\gamma_{\mathrm{PD}}$ is the pure‑dephasing rate, and $\gamma_{\mathrm{AD}}$ is the amplitude‑damping rate. Here, we define
$\mathcal{L}_{0}$ $\coloneqq\frac{i\omega_{z}}{2}[\sigma_{z},\rho]+\gamma_{\mathrm{AD}}\!\Bigl(\sigma_{-}\rho\sigma_{+}-\tfrac{1}{2}\{\sigma_{+}\sigma_{-},\rho\}\Bigr)$
$\mathcal{L}_{1}$ $\coloneqq\gamma_{\text{PD}}\bigr(\sigma_{z}\rho\sigma_{z}-\rho\bigr).$ (18)
To construct the memory kernel of the (PMME), it is convenient to diagonalise Eq. (III.4) in a bi‑orthonormal set of right/left eigenoperators (where we call “damped basis”),
$\mathcal{L}(R_{i})=\lambda_{i}R_{i},\quad\mathcal{L}^{\dagger}(L_{i})=\lambda_{i}L_{i},\quad\mathrm{Tr}\!\bigl[L_{i}R_{j}\bigr]=\delta_{ij}.$ (19)
Note that in our case, $\mathcal{L}_{0},\mathcal{L}_{1}]$, therefore, these two operator are simultaneusly diagonalizable with eigenvalue $\lambda^{0}_{i},\lambda^{1}_{i}$.
By diagonalizing the superoperator in Eq. (III.4), the corresponding eigenoperators and eigenvalus are given by
$\left\{\begin{array}[]{lll}R_{0}=\frac{I+\sigma_{z}}{\sqrt{2}},&L_{0}=\frac{I}{\sqrt{2}},&\lambda_{0}=0\\ R_{1}=\frac{\sigma_{z}}{\sqrt{2}},&L_{1}=\frac{\sigma_{z}-I}{\sqrt{2}},&\lambda_{1}=-\gamma_{\text{AD}}\\ R_{2}=\sigma_{-},&L_{2}=\sigma_{+},&\lambda_{2}=i\omega_{z}-(2\gamma_{\text{PD}}+\frac{1}{2}\gamma_{\text{AD}})\\ R_{3}=\sigma_{+},&L_{3}=\sigma_{-},&\lambda_{3}=-i\omega_{z}-(2\gamma_{\text{PD}}+\frac{1}{2}\gamma_{\text{AD}})\end{array}\right$. (24)
where for $\mathcal{L}_{0}$ and $\mathcal{L}_{1}$, the eigenvalues are
$\left\{\begin{array}[]{ll}\lambda^{0}_{0}=0,&\lambda^{1}_{0}=0\\ \lambda^{0}_{1}=-\gamma_{\text{AD}},&\lambda_{1}^{1}=0\\ \lambda^{0}_{2}=i\omega_{z}-\frac{1}{2}\gamma_{\text{AD}},&\lambda^{1}_{2}=-2\gamma_{\text{PD}}\\ \lambda_{3}^{0}=-i\omega_{z}-\frac{1}{2}\gamma_{\text{AD}},&\lambda^{1}_{3}=-2\gamma_{\text{PD}}\end{array}\right$. (29)
Any state can be expanded in this “damped basis” as
$\rho(t)=\sum_{i=0}^{3}\mu_{i}(t)R_{i}$ (30)
Inserting the expansion into the post‑Markovian master equation gives the scalar form
$\dot{\mu}_{i}(t)=\lambda_{i}^{0}\mu_{i}(t)+\lambda_{i}^{1}\int_{0}^{t}d\tau e^{\lambda_{i}\tau}k(\tau)\mu_{i}(t-\tau)$ (31)
by taking the Laplace transform to the $s$-domain we can formally solve the linear combination coefficient
$\tilde{\mu}_{i}(s)=\frac{\mu_{i}(0)}{s-\lambda_{i}^{0}-\lambda^{1}_{i}\tilde{k}(s-\lambda_{i})}$ (32)
Here $\tilde{k}(s)$ and $\hat{\mu}_{i}(s)$ are the Laplace transforms of the memory kernel $k(t)$ and of $\mu_{i}(t)$, respectively. In the Markovian limit of $k(t)=\delta(t)$, one recovers $\dot{\mu}_{i}(t)=\lambda_{i}\mu_{i}(t)$ and thus Eq. (III.4).
This diagonal form greatly simplifies both analytical manipulations and the numerical extraction of $\mu_{i}(t)$ from tomography data on the IBM superconducting processor.
IV Experiment Setup
In this section, we describe our experimental setup and results for probing the non-Markovianity of the noise process introduced in the previous section. We select a four-qubit subsystem on an IBM superconducting processor, designating one “main” qubit whose dynamics we monitor, and three spectator qubits that serve as a controlled environment. By initializing all three spectators in the same state, we isolate and study their collective effect on the main qubit’s evolution.
IV.1 Quantum State Tomography
To track the time-dependent density matrix $\rho(t)$ of the main qubit, we perform quantum state tomography at each probe time $t$. More concretely, we choose the tomographically complete POVM $\{E_{m,k}\}$ such that each outcome and corresponding probability is
Figure 3: Layout of the IBM superconducting quantum processor in the heavy-hex lattice (partial view). Each qubit connects to at most three neighbors. In this example, we select qubits 23, 24, 25, and 34, with qubit 24 serving as the main qubit and the other three acting as spectator qubits.
$p_{m,k}=\mathrm{Tr}(\rho E_{m,k}).$ (33)
where $m$ labels the measurement basis while $k$ labels the outcome. For example, in the single qubit case, the $m$ labels the pauli operators $\sigma_{x},\sigma_{y},\sigma_{z}$ while the $k$ labels the outcome of getting 0 or 11. We will repeat the measurement circuit with sufficiently large number of shots to ensure that the empirical estimation $\hat{p}_{m,k}$ is good enough to approximate the true probability $p_{m,k}$.
Figure 4: Schematic of the basic experimental setup. The circuit is composed of three parts: (1) the state‐preparation unitary $U_{\mathrm{SP}}$, (2) an idling phase implemented by a sequence of identity gates, and (3) the final basis rotations for tomography. By varying the number of identity gates, we control how long the main qubit idles and interacts with its environment. After collecting the measurement statistics, we form a data vector of empirical probabilities and the corresponding POVM operators for either state or process tomography and solve the resulting least‐squares problem. The raw estimates of the density matrix $\hat{\rho}$ and the Choi matrix $\hat{\chi}_{t}$ are then projected onto the physical subspace to enforce the required properties of a valid quantum state or a trace‐preserving, completely positive map.
All of the estimated probabilities $\hat{p}_{m,k}(t)$ (here we use $t$ to label the time point $t$) can be stacked into a vector $y(t)$. We can also provide that $x(t)$ denotes the vectorized density operator $|\rho(t)\rangle\!\rangle$. Now, we can formulate the reconstruction problem into solving a system of linear equations:
$A\,x(t)\approx y(t),\quad A_{(m,k),\alpha\beta}=\bigl\langle\!\langle\alpha\beta\big|E_{m,k}\bigr\rangle\!\bigr\rangle,$ (34)
where each row of AA is the vectorized POVM element $|E_{m,k}\rangle\!\rangle$, and the matrix multiplication of AA and $x(t)$ is equivalent to applying $\mathrm{Tr}(\rho E_{m,k})$. We first compute the raw least-squares estimate
$\hat{x}(t)=A^{+}\,y(t)\quad\Longrightarrow\quad\hat{\rho}(t),$ (35)
where $A^{+}$ denotes the pseudo inverse. Since the estimated density matrix $\hat{\rho}$ will not always be positive and with unit trace, we need to enforce physicality such as Hermiticity, positive semidefiniteness, and unit trace by either
  1. 1.
    projecting any negative eigenvalues of $\hat{\rho}(t)$ to zero and renormalizing, or
  2. 2.
    solving the constrained optimization
    $\min_{\rho\succeq 0,\;\mathrm{Tr}\rho=1}\sum_{m,k}\bigl[\mathrm{Tr}(E_{m,k}\rho)-\hat{p}_{m,k}(t)\bigr]^{2},$
to ensure the final $\hat{\rho}(t)$ is a valid density matrix that best fits our observed data.
IV.2 Quantum Process Tomography
Quantum process tomography is an extension of the methods for quantum state tomography characterizing the unknown quantum channel $\mathcal{E}$ rather than a fixed state $\rho$. To do process tomography, we need to both prepare the input state and measure the output in different bases to fully reconstruct the process matrix (or Choi matrix). In practice, we prepare a tomographically complete set of input states $\{\rho_{p}\}$. For example, the initial states are commonly chosen as $\{|0\rangle,|1\rangle,|+\rangle,|+i\rangle\}$ for the single qubit process tomography. Similar to quantum state tomography, we measure the output state in a complete set of POVM $\{E_{m,k}\}$ and denote the measurement probability as
$\hat{p}_{p,m,k}=\mathrm{Tr}(E_{m,k}\mathcal{E}_{t}(\rho_{p}))$ (36)
where $p$ labels the initial state. Note that we can relate this probability to the Choi matrix by the Choi Jamilkowski isomorphism
$\hat{p}_{p,m,k}=\mathrm{Tr}(\chi_{t}(\rho_{p}^{T}\otimes E_{m,k}))$ (37)
where $\chi_{t}$ is Choi matrix of the quantum channel $\mathcal{E}_{t}$. We can, again, vectorize each quantity for algebraic convenience by stacking the outcome $\hat{p}_{p,m,k}$ into a vector $y$ and also defining
$x(t)=|\chi_{t}\rangle\!\rangle,\quad A_{(p,m,k),\alpha\beta}=\langle\!\langle\alpha\beta|\rho_{p}^{T}\otimes E_{m,k}\rangle\!\rangle.$ (38)
Therefore, the tomography reconstruction can again become solving a system of linear equation
$Ax(t)\approx y(t).$ (39)
By solving the least square problem, we have the raw estimation $\hat{x}(t)=A^{+}y(t)$. In order to make the estimated data physical (a CPTP map), we project the resulting Choi matrix $\chi_{t}$ such that $\chi_{t}\geq 0$ and $\mathrm{Tr}_{1}\chi_{t}=I$. An alternative way to deal with the post processing is to solve the constraint optimization problem
$\min_{\chi_{t}\geq 0,\mathrm{Tr}_{1}\chi_{t}=I}\sum_{p,m,k}\left(\mathrm{Tr}(\chi_{t}(\rho_{p}^{T}\otimes E_{m,k}))-\hat{p}_{p,m,k}\right)^{2}.$ (40)
to obtain a Choi matrix $\chi_{t}$ associated to a CPTP map and also closest to the raw estimation probability $\hat{p}_{p,m,k}$.
IV.3 Memory Kernel Reconstruction
To reconstruct the memory kernel $k(t)$ from the tomographically estimated density matrices $\hat{\rho}(t)$, we start from the Laplace‐domain relation in Eq. (32). For convenience, we define the normalized functions
$\xi_{i}(t)\coloneqq\frac{\mu_{i}(t)}{\mu_{i}(0)}.$ (41)
Rewriting Eq. (32) then gives
$\tilde{k}\bigl(s-\lambda_{i}\bigr)$ $=\frac{1}{\lambda_{i}^{1}}\Bigl[s-\lambda_{i}^{0}-\tfrac{1}{\tilde{\xi}_{i}(s)}\Bigr].$ (42)
Applying the inverse Laplace transform yields
$k(t)$ $=\frac{e^{-\lambda_{i}t}}{\lambda_{i}^{1}}\,\text{Lap}^{-1}\!\Bigl[s-\lambda_{i}^{0}-\tfrac{1}{\tilde{\xi}_{i}(s)}\Bigr],$ (43)
which allows us to extract $k(t)$ once $\xi_{i}(t)$ is known.
This expression is valid only if $\lambda_{i}^{1}\neq 0$. From the eigenvalue analysis in Section III.4, we have
$\lambda_{0}^{1}=\lambda_{1}^{1}=0,\quad\lambda_{2}^{1}=\lambda_{3}^{1}=-2\gamma_{\mathrm{PD}}.$ (44)
Hence, only the modes $i=2,3$ yield nontrivial memory‐kernel contributions. In that subspace the right eigenoperators are
$R_{2}=\sigma_{-},\quad R_{3}=\sigma_{+}=R_{2}^{\dagger}.$ (45)
By Hermiticity of the density matrix ($\rho^{\dagger}=\rho$), it follows immediately that
$\xi_{2}(t)=\xi_{3}^{*}(t),$ (46)
so the two coefficients are complex conjugates rather than independent functions. Now, if we define
$k_{i}(t)=\frac{e^{-\lambda_{i}t}}{\lambda_{i}^{1}}\,\text{Lap}^{-1}\!\Bigl[s-\lambda_{i}^{0}-\tfrac{1}{\tilde{\xi}_{i}(s)}\Bigr],$ (47)
one infers from Eq. (46) that
$k_{2}(t)=k_{3}^{*}(t).$ (48)
A detailed derivation can be found in Appendix B. Consequently, if $k(t)$ is treated as a scalar, the two reconstructions generally differ, $k_{2}(t)\neq k_{3}(t)$.
To resolve this ambiguity, we can lift the memory kernel to a superoperator
$k(\tau)X=\sum_{i=0}^{3}k_{i}(\tau)\,R_{i}\,\mathrm{Tr}(L_{i}X),$ (49)
so that its action on the delayed state $\rho(t-\tau)$ reads
$k(\tau)\rho(t-\tau)=\sum_{i=0}^{3}k_{i}(\tau)\,\mu_{i}(t-\tau)\,R_{i}.$ (50)
In this formulation each mode has its own kernel function $k_{i}(\tau)$; if all $k_{i}(\tau)$ coincide, the superoperator reduces to the scalar memory kernel.
Now, we discuss how to reconstruct the empirical $k_{i}(\tau)$ from the tomography data.
  1. 1.
    We uniformly sample the time points $0,T_{\max}]$ with time step $\Delta t$ and total number of points $N=T_{\max}/\Delta t+1$. For each time point, we do the quantum state tomography to get the estimated quantum state $\hat{\rho}(t)$.
  2. 2.
    We calculate the coefficient $\mu_{i}(t)$ by $\mu_{i}(t)=\mathrm{Tr}(L_{i}\hat{\rho}(t))$ and $\xi_{i}(t)=\mu_{i}(t)/\mu_{i}(0)$.
  3. 3.
    For the mode with $\lambda_{i}^{1}\neq 0$, we do the numerical Laplace transform to get $\tilde{\xi}_{i}(s)$.
  4. 4.
    We perform the numerical inverse Laplace transform on $\frac{1}{\lambda_{i}^{1}}\left[s-\lambda_{i}^{0}-\frac{1}{\tilde{\xi}_{i}(s)}\right]$ and multiply the resulting function by $e^{-\lambda_{i}t}$.
Below, we discuss how to do the numerical Laplace and inverse Laplace transform.
IV.3.1 Numerical Laplace Transform
Let $f(t)$ be a causal function ($f(t)=0$ for $t<0$). We sample $f(t)$ on a uniform time grid
$t_{n}=n\Delta t,\qquad n=0,\dots,N-1,\qquad\Delta t=\frac{T_{\max}}{N-1}.$ (51)
Because the data have are finite in duration, we approximate the Laplace integral with an upper limit $T_{\max}$:
$\tilde{f}(\sigma+i\omega)\;\approx\;\int_{0}^{T_{\max}}f(t)\,e^{-\sigma t}\,e^{-i\omega t}\,dt.$ (52)
With only a finite set of samples, the integral is further approximated by a Riemann sum, yielding a discrete‐time Fourier transform (DTFT) in damped form:
$\tilde{f}(\sigma+i\omega_{k})\;\approx\;\Delta t\sum_{n=0}^{N-1}f(t_{n})\,e^{-\sigma t_{n}}\,e^{-i\omega_{k}t_{n}},$ (53)
where
$\omega_{k}=\frac{2\pi k}{N\Delta t},\qquad k=0,\dots,N-1.$ (54)
In practice we set the damping parameter to $\sigma=\tfrac{1}{3T_{\max}}$.
IV.3.2 Numerical Inverse Laplace Transform
The inverse Laplace transform can be evaluated via the Bromwich integral,
$f(t)=\frac{1}{2\pi}\int_{-\infty}^{\infty}\tilde{f}(\sigma+i\omega)\,e^{\sigma t}\,e^{i\omega t}\,d\omega,$ (55)
which we approximate by again truncating the frequency range and discretising:
$f(t_{n})$ $=\frac{e^{\sigma t_{n}}}{2\pi}\,\Delta\omega\sum_{k=-N/2}^{N/2-1}\tilde{f}(\sigma+i\omega_{k})\,e^{i\omega_{k}t_{n}},$ (56)
with $\Delta\omega=\tfrac{2\pi}{N\Delta t}.$ Because the sum is periodic, we may shift the index range to $k=0,\dots,N-1$ without changing the value. The Bromwich integral thus reduces to an inverse discrete Fourier transform:
$f(t_{n})=\frac{e^{\sigma t_{n}}}{N\Delta t}\sum_{k=0}^{N-1}\tilde{f}(\sigma+i\omega_{k})\,e^{i\omega_{k}t_{n}}.$ (57)
V Results
In this section we apply a suite of information–theoretic diagnostics to the experimentally reconstructed dynamical map $\Phi_{t}$. These tests probe both crosstalk between qubits and memory‐induced non‐Markovian effects. Specifically, we will (i) test CP-divisibility via the Choi matrix, (ii) examine information backflow through trace distance and relative entropy, (iii) quantify crosstalk via two-qubit quantum mutual information, and (iv) reconstruct the PMME memory kernel.
V.1 CP–divisibility test
As discussed in Sec. III.1, a dynamical map $\Phi_{t}$ is CP–divisible if every intermediate propagator
$\mathcal{V}_{t,s}$ $=\Phi_{t}\circ\Phi_{s}^{-1}$ (58)
is completely positive for all $0\leq s\leq t$. Equivalently, the Choi matrix $\chi_{t,s}$ of $\mathcal{V}_{t,s}$ must be positive semidefinite. In practice, we proceed as follows:
  1. 1.
    Reconstruct the Choi matrix $\chi_{t}$ at each time $t$ by process tomography.
  2. 2.
    Reshuffle $\chi_{t}$ into the superoperator matrix $S_{t}=\mathcal{R}(\chi_{t})$.
  3. 3.
    Compute the intermediate superoperator matrix, since $\Phi_{s}$ may not always invertible, we apply the pseudo-inverse on the superoperator matrix
    $S_{t,s}$ $=S_{t}\,S_{s}^{+}\,$, (59)
    and recover its Choi matrix
    $\chi_{t,s}$ $=\mathcal{R}^{-1}\bigl(S_{t,s}\bigr)=\mathcal{R}^{-1}\bigl(S_{t}\,S_{s}^{+}\bigr).$ (60)
  4. 4.
    Evaluate the minimum eigenvalue
    $\lambda_{\min}\bigl[\chi_{t,s}\bigr]$ $\geq 0\quad\Longleftrightarrow\quad\mathcal{V}_{t,s}\ \text{is CP.}$ (61)
Figure 5 shows a triangular heat–map of $\lambda_{\min}[\chi_{t,s}]$ over the domain $0\leq s\leq t\leq T$. We use the “magma” colormap such that for the light yellow region, $\lambda_{\min}\approx 0$ which indicates CP–preserving regions. While for the orange to dark red region $\lambda_{\min}<0$ and purple $\lambda_{\min}\ll 0$ mark violations of CP–divisibility. The presence of extended red–purple patches in the plots provide clear evidence of non–Markovian memory effects in our system.
Figure 5: Heat‐maps of the minimum eigenvalue $\lambda_{\min}\bigl[\chi_{t,s}\bigr]$ for the intermediate Choi matrix as a function of start time $s$ and idle time $t$, measured on IBM’s 127-qubit “Strasbourg” processor. (a) Left panel shows the case where the main qubit is number 4 and the spectator qubits are {3, 5, 15}; (b) Right panel uses main qubit 25 with spectators {26, 27, 16}. In each experiment the spectator qubits are initialized in the $|+\rangle$ state and held idle for duration $s$, after which full process tomography is performed on the main qubit at time $t$. The triangular region $s>t$ is masked. Regions with $\lambda_{\min}\approx 0$ (light yellow) are CP‐divisible, whereas negative values (orange to purple) indicate CP‐divisibility violations (non-Markovianity).
V.2 Information Backflow
A second signature of non-Markovianity is the backflow of the information from environment to system. As we discussed in the Section. III.2, both trace distance and quantum relative entropy are contractive under a CPTP map $\Phi_{t}$:
$T(\Phi_{t}(\rho_{1}),\Phi_{t}(\rho_{2}))\leq T(\rho_{1},\rho_{2})$ (62)
$D(\Phi(\rho_{1})\|\Phi(\rho_{2}))\leq D(\rho_{1}\|\rho_{2})$ (63)
For Markovian noise, the map is CP-divisible. Therefore, these two quantities are monotonically decreasing over time in the Markovian case, and the increase of these two quantities signifies that the information flows back from the environment to the system.
In our experiments, we consider two orthogonal input pairs:
$\rho_{1}(0)=|+\rangle\langle+|,\quad$ $\rho_{2}(0)=|-\rangle\langle-|,$ (64)
$\rho_{1}(0)=|+i\rangle\langle+i|,\quad$ $\rho_{2}(0)=|-i\rangle\langle-i|,$ (65)
with spectator qubits initialized either in $|+\rangle^{\otimes 3}$ or $|1\rangle^{\otimes 3}$. At each idle time $t$ we perform full state tomography on the main qubit, reconstruct $\rho_{1,2}(t)$, and compute both $T\bigl(\rho_{1}(t),\rho_{2}(t)\bigr)$ and $D\bigl(\rho_{1}(t)\|\rho_{2}(t)\bigr)$.
Fig. 6 shows the time evolution of the trace distance $T(\rho_{1}(t),\rho_{2}(t))$ and quantum relative entropy $D(\rho(t_{1})\|\rho(t_{2}))$. The periodic revival of these two quantities demonstrates non‐Markovian information backflow from the environment into the system. At the same time, the overall amplitude of these oscillations is enveloped by an exponential decay function, reflecting the underlying amplitude‐damping effect induced by spontaneous photon emission. In the long‐time limit all trajectories relax to the ground state $|0\rangle$, where both $T$ and $D$ asymptotically vanish.
Figure 6: Time evolution of (a) the trace distance $T\bigl(\rho_{1}(t),\rho_{2}(t)\bigr)$ and (b) the quantum relative entropy $D\bigl(\rho_{1}(t)\,\|\,\rho_{2}(t)\bigr)$ for the two initial pairs $(|+\rangle,|-\rangle)$ and $(|+i\rangle,|-i\rangle)$. The experiments are conducted on IBM’s 127-qubit “Brussels” processor, the main qubit index is 49 while the spectator qubits index are {48, 50, 55}. Solid lines correspond to spectator qubits in $|+\rangle^{\otimes 3}$; dashed lines to spectators in $|1\rangle^{\otimes 3}$. The periodic revivals indicate information backflow and thus non-Markovian behavior.
V.3 Information‑theoretic characterization of qubit crosstalk
A convenient way to quantify correlations generated between two idle qubits AA and $B$ is the quantum mutual information
$I(A\!\mathrel{\mathop{\ordinarycolon}}\!B)_{\rho(t)}$ $=S(A)_{\rho(t)}+S(B)_{\rho(t)}-S(AB)_{\rho(t)}$
$=D\bigl(\rho_{AB}(t)\bigl\|\rho_{A}(t)\!\otimes\rho_{B}(t)\bigr),$ (66)
where $D(\cdot\|\cdot)$ is the quantum relative entropy. Because $D(\rho\|\sigma)=0$ if and only if $\rho=\sigma$, a non‑zero value of $I(A\!\mathrel{\mathop{\ordinarycolon}}\!B)$ implies that $\rho_{AB}(t)$ is not a product state, i.e. some qubit–qubit correlation has been established.
A complementary indicator is the quantum conditional entropy
$S(A|B)_{\rho(t)}$ $=S(AB)_{\rho(t)}-S(B)_{\rho(t)}$
$=-\,D\!\bigl(\rho_{AB}(t)\,\bigl\|\,I_{A}\!\otimes\!\rho_{B}(t)\bigr),$ (67)
which has the following upper and lower bound
$-\log|A|\leq S(A|B)_{\rho(t)}\leq\log|A|.$ (68)
For two qubits the range is $1\leq S(A|B)\leq 1$. The lower bound for $S(A|B)_{\rho(t)}$ saturates for the maximally entangled state $|\Phi^{+}\rangle_{AB}$, and the upper bound saturates for the product state $\Pi_{A}\otimes\sigma_{B}$ where $\Pi_{A}=\frac{1}{|A|}I_{A}$ (maximally mixed state for system AA) and $\sigma_{B}$ is any valid quantum state for system $B$.
A critical property for quantum conditional entropy is $S(A|B)_{\rho(t)}\geq 0$ if $\rho_{AB}(t)$ is separable. Therefore, a negative conditional entropy is a sufficient witness of entanglement ($S(A|B)<0\!\;\Rightarrow\;$ $\rho_{AB}$ is non‑separable), whereas $S(A|B)>0$ only rules out this particular witness and is compatible with purely classical or weakly quantum correlations.
We performed four independent runs on IBM’s 127‑qubit Sherbrooke processor, each starting from an orthogonal product state
$\rho_{AB}(0)$ $=|+\rangle\langle+|^{\otimes 2},$ $\rho_{AB}(0)$ $=|-\rangle\langle-|^{\otimes 2},$
$\rho_{AB}(0)$ $=|+i\rangle\langle+i|^{\otimes 2},$ $\rho_{AB}(0)$ $=|-i\rangle\langle-i|^{\otimes 2}.$ (69)
Both qubits are left idle for a variable duration $t$. At each time point we carry out full two‑qubit state tomography, obtain $\rho_{AB}(t)$ and its marginals, and compute $I(A\!\mathrel{\mathop{\ordinarycolon}}\!B)_{\rho(t)}$ as well as $S(A|B)_{\rho(t)}$. For qubits, the mutual information is bounded by  $0\leq I(A\mathrel{\mathop{\ordinarycolon}}B)_{\rho(t)}\leq 2$. Fig. 7(a) shows that $I(A\mathrel{\mathop{\ordinarycolon}}B)_{\rho(t)}$ quickly departs from zero, peaks at $t\!\approx\!6\;\mu\mathrm{s}$, and exhibits pronounced revivals at $\sim 30$, $60$, and $90\;\mu\mathrm{s}$—closely matching the backflow periods found earlier. In Fig. 7(b) these revivals correspond to dips of $S(A|B)$ below zero, confirming that genuine entanglement is generated during the information exchange between two qubits. The overall envelope of both curves decays due to amplitude damping, eventually driving the two‑qubit state towards $\lvert 0\rangle^{\otimes 2}$, where mutual information and conditional entropy both vanish.
Figure 7: (a) Quantum mutual information $I(A\!\mathrel{\mathop{\ordinarycolon}}\!B)_{\rho(t)}$ and (b) quantum conditional entropy $S(A|B)_{\rho(t)}$ for four different initial product states. An initial rise of $I(A\!\mathrel{\mathop{\ordinarycolon}}\!B)$ up to $\sim 0.7$ bits within the first $10\;\mu\mathrm{s}$ demonstrates fast crosstalk‑induced correlation build‑up. Revivals every ${\sim}30\;\mu\mathrm{s}$ coincide with the non‑Markovian signatures observed in Secs. III.1III.2. The conditional entropy remains negative for the strongest revival peaks, certifying transient entanglement between the qubits, while it relaxes towards zero as amplitude damping drives both qubits to $\lvert 0\rangle$.
V.4 Reconstruction of the PMME memory kernel
To characterise the environment’s memory effects, we reconstruct the memory kernel $k(t)$ of the post‑Markovian master equation
$\dot{\rho}(t)$ $=\mathcal{L}_{0}\,\rho(t)+\int_{0}^{t}\!d\tau\;e^{(\mathcal{L}_{0}+\mathcal{L}_{1})\tau}\,\mathcal{L}_{1}\,k(\tau)\,\rho(t-\tau),$ (70)
where $\mathcal{L}_{0}$ is a best‑fit Lindblad generator obtained by parameter optimisation and $\mathcal{L}_{1}$ accounts for non‑Markovian effects; in our model $\mathcal{L}_{1}$ contains only a dephasing term. The corresponding Lindblad equation reads
$\dot{\rho}\;=\;(\mathcal{L}_{0}+\mathcal{L}_{1})\rho.$ (71)
We conduct our experiment and kerne reconstruction in the following procedure:
  1. 1.
    Tomography and Bloch‑vector extraction. For the four initial states $|\pm\rangle,\,|\pm i\rangle$ we perform state tomography at $50$ time points. At each point we append $50$ identity gates (gate duration $\sim 56\,$ns) and record the Bloch components $v_{x,y,z}(t)$ together with the purity ${\rm Tr}\rho^{2}(t)$.
  2. 2.
    Lindblad trajectory. We fit a Lindblad model to the trajectories $v_{i}(t)$ by minimizing the mean‑squared error
    $\mathrm{MSE}=\frac{1}{3N}\sum_{i\in\{x,y,z\}}\sum_{n=1}^{N}\bigl|v_{i}(t_{n})-\hat{v}_{i}(t_{n})\bigr|^{2},$ (72)
    where $\hat{v}_{i}(t_{n})$ denotes the Bloch components generated by the Lindblad equation. A gradient‑based optimizer is used to determine the parameters $\omega_{z}$, $\gamma_{\mathrm{AD}}$, and $\gamma_{\mathrm{PD}}$ (see Eq. (III.4)) until the loss converges.
  3. 3.
    Memory‑kernel reconstruction. Following the procedure in Section. IV.3, we perform a numerical Laplace transform to obtain the coefficients $\tilde{\xi}(\sigma+i\omega_{k})$ for $k=0,\dots,N-1$. Substituting these coefficients into Eq. (42) and applying a numerical inverse Laplace transform yield the reconstructed kernel $k_{i}(t)$.
Figure 8: Reconstruction of the post‑Markovian memory kernel for four different initial states of the main qubit. Each column corresponds to a different preparation, indicated above the panels. (Rows 1–3) Masured Bloch‑vector components $v_{x}(t)$ (yellow), $v_{y}(t)$ (orange), and $v_{z}(t)$ (red). Markers are experimental data; solid lines are the best‑fit Markovian (Lindblad) trajectories obtained from the MSE error minimization as described in the text. The Lindblad model reproduces the overall exponential damping but fails to capture the pronounced revivals at $t\approx 25$, $60$, and $95,\mu\mathrm{s}$, signalling non‑Markovian backflow. (Row 4) Purity ${\rm Tr}\bigl[\rho(t)^{2}\bigr]$ for the same data sets (markers) and the corresponding Markovian prediction (solid green). The oscillations about the Markovian envelope again reflect the memory‑induced deviations. (Row 5) Reconstructed memory kernel $k_{3}(t)$ obtained from the inversion procedure (Section.IV.3); real part (blue circles) and imaginary part (pink triangles) are shown. All four initial conditions give nearly identical kernels: a damped oscillation at $\omega/2\pi\approx 13,$kHz, confirming that the non‑Markovianity is an intrinsic property of the device rather than an artifact of state preparation.
Fig. 8 (rows 1–4) compares the experimental Bloch coordinates (markers) with the purely Markovian prediction (solid lines). Although $\mathcal{L}_{0}$ reproduces the overall damped‑oscillation trend, it fails to capture the revivals at $t\!\approx\!25,\,60,$ and $95\,\mu\mathrm{s}$—the same times at which information back‑flow was observed earlier. These non‑Markovian effects are most pronounced in the $v_{x}$ and $v_{y}$ components, whereas $v_{z}$ shows an almost exponential decay. Consequently, the Markovian part can be identified with the in‑plane rotation term $i\omega_{z}[\sigma_{z},\rho]$ and the amplitude‑damping contribution $\gamma_{\mathrm{AD}}\bigl(\sigma_{-}\rho\sigma_{+}-\tfrac{1}{2}\{\sigma_{+}\sigma_{-},\rho\}\bigr)$. The purity predicted by the Lindblad model clearly displays an initial exponential decrease (pure dephasing) followed by a late‑time increase (amplitude damping). The experimental purity oscillates about this curve, indicating a combination of long‑time memoryless effects and short‑time non‑Markovian contributions.
The reconstructed kernel $k_{3}(\tau)$, shown in the bottom row, exhibits a damped oscillation with frequency $\omega/2\pi\approx 13\,\mathrm{kHz}$. Remarkably, the kernel is state‑independent: all four initial conditions yield almost identical $k_{3}(\tau)$, confirming that the observed non‑Markovianity is an intrinsic property of the device rather than a state‑preparation artifact. The strong deviation of $k_{3}(\tau)$ from a delta function highlights the significant memory effect and continual exchange of information between system and environment.
VI Conclusion
[H] In this work, we presented an information-theoretic, experiment‑driven characterization of non‑Markovian noise on IBM superconducting qubits. By combining process tomography, information‑theoretic witnesses and an explicit reconstruction of the Post‑Markovian Master Equation (PMME) kernel, we obtained a self‑consistent picture of how memory and crosstalk shape single‑qubit dynamics on a realistic 127‑qubit processor. The CP-divisibility test revealed large triangular regions in the $s,t$ plane where the intermediate map failed to remain completely positive, demonstrating a clear breakdown of CP‑divisibility. Also in the information backflow characterization of the non-Markovianity, we identified pronounced revivals in both the trace distance and the quantum relative entropy of initially orthogonal states, confirming that information lost to the environment flows back into the qubit at regular intervals. Direct two‑qubit probes showed that these revivals coincide with the rapid growth of quantum mutual information and transiently negative conditional entropy, establishing that short‑range qubit–qubit crosstalk is the dominant physical origin of the observed memory effects.
Starting from a Lindblad approximation that faithfully reproduces the long‑time exponential damping, we analytically calculated the PMME kernel in Laplace space and reconstructed a nearly state‑independent memory kernel. This memory kernel has oscillating behavior in both the real and imaginary parts which generally deviates from the delta function kernel that results in the Markovian effect, thereby validating the kernel extraction procedure and providing a quantitative reduced model of the device noise. Taken together, these results show that present superconducting devices possess hardware‑level memory that is not fully characterized by the standard $T_{1}/T_{2}$ time and must be addressed to reach fault‑tolerant thresholds.
Appendix A relation between superoperator matrix and Choi matrix
In this section we derive the relation between the superoperator matrix $S_{t}$ and the Choi matrix $\chi_{t}$. Recalled that by Choi–Jamiołkowski isomorphism we have
$\Phi_{t}(\rho)$ $=\mathrm{Tr}_{1}\bigl[\chi_{t}(\rho^{T}\otimes I)\bigr]$
$=\sum_{k}(\langle k|\otimes I)\chi_{t}(\rho^{T}\otimes I)(|k\rangle\otimes I)$
$=\sum_{k,\ell}(\langle k|\otimes I)\chi_{t}(|\ell\rangle\langle\ell|\otimes I)(\rho^{T}\otimes I)(|k\rangle\otimes I)$
$=\sum_{k,\ell}(\langle k|\otimes I)\chi_{t}(|\ell\rangle\otimes I)\rho_{k\ell}$
$=\sum_{ijk\ell}(\chi_{t})_{ki,\ell j}\rho_{k\ell}|i\rangle\langle j|$ (73)
Therefore,
$[\Phi_{t}(\rho)]_{ij}=\sum_{k\ell}(\chi_{t})_{ki,\ell j}\rho_{k\ell}$ (74)
and the vectorized map of a density matrix $\rho$ is defined by stacking the column of $\rho$, therefore, $\text{vec}(\rho)]_{ij}=\rho_{ij}$. By comparing the equation above with the vectorized superoperator
$[\text{vec}(\Phi_{t}(\rho))]_{ij}$ $=\sum_{k\ell}(S_{t})_{ij,k\ell}[\text{vec}(\rho)]_{k\ell}$ (75)
we immediately see that
$(S_{t})_{ij,k\ell}=(\chi_{t})_{ki,\ell j}$ (77)
while the inverse map is
$(\chi_{t})_{ij,k\ell}=(S_{t})_{j\ell,ik}$ (78)
as stated in the main text.
Appendix B complex conjugation relation between $k_{2}(t)$ and $k_{3}(t)$
In this section we derive the relation $k_{2}(t)=k_{3}^{*}(t)$. Recall that the Laplace transform of the complex conjugate of a function $g(t)$ satisfies
$\text{Lap}\bigl[g^{*}(t)\bigr]$ $=\int_{0}^{\infty}g^{*}(t)\,e^{-st}\,dt$ (79)
$=\Bigl(\int_{0}^{\infty}g(t)\,e^{-s^{*}t}\,dt\Bigr)^{*}$ (80)
$=\tilde{g}^{*}\bigl(s^{*}\bigr).$ (81)
Starting from the Laplace‐domain expression for $k_{2}(t)\,e^{\lambda_{2}t}$,
$\tilde{k}_{2}\bigl(s-\lambda_{2}\bigr)$ $=\frac{1}{\lambda_{2}^{1}}\Bigl[s-\lambda_{2}^{0}-\tfrac{1}{\tilde{\xi}_{2}(s)}\Bigr],$ (82)
we substitute $s\to s^{*}$ and take the complex conjugate. This gives the Laplace transform of $\bigl[k_{2}(t)e^{\lambda_{2}t}\bigr]^{*}$:
$\tilde{k}_{2}^{*}\bigl(s^{*}-\lambda_{2}\bigr)$ $=\frac{1}{\lambda_{2}^{1*}}\Bigl[s-\lambda_{2}^{0*}-\tfrac{1}{\tilde{\xi}_{2}^{*}(s^{*})}\Bigr].$ (83)
Using $\lambda_{2}^{1}=\lambda_{3}^{1}=-2\gamma_{\mathrm{PD}}$, $\lambda_{2}^{0*}=\lambda_{3}^{0}$, and $\tilde{\xi}_{2}^{*}(s^{*})=\tilde{\xi}_{3}(s)$, we recognize the right‐hand side as
$\frac{1}{\lambda_{3}^{1}}\Bigl[s-\lambda_{3}^{0}-\tfrac{1}{\tilde{\xi}_{3}(s)}\Bigr]$ $=\text{Lap}\bigl[k_{3}(t)\,e^{\lambda_{3}t}\bigr].$ (84)
Therefore,
$\bigl[k_{2}(t)\,e^{\lambda_{2}t}\bigr]^{*}$ $=k_{3}(t)\,e^{\lambda_{3}t}\;\Longrightarrow\;k_{2}^{*}(t)=k_{3}(t),$ (85)
as claimed.
Appendix C $Z$ crosstalk effect
This section investigates how ZZ–type interactions modify the dynamics of a designated main qubit. Our analysis progresses through three levels of complexity:
  1. (i)
    Pure $Z$ coupling. We start with the minimal coherent model in which the main qubit interacts with a single spectator qubit through a $Z$ Hamiltonian. No decoherence is included at this stage.
  2. (ii)
    $Z$ coupling with local noise. Next, we incorporate independent dephasing and amplitude–damping channels on both qubits to contrast purely coherent and open-system dynamics.
  3. (iii)
    Multi-spectator scenario with global dissipation. Finally, we generalise to a realistic device layout where the main qubit experiences $Z$ crosstalk with multiple spectators, and every qubit is simultaneously subject to dephasing and amplitude damping.
At each step we write down the corresponding Lindbladian, solve the Heisenberg-picture equations of motion for the relevant operator set, and highlight the resulting modifications to the Bloch-vector trajectories of the main qubit.
C.1 $Z$ crosstalk between two qubits
The simplest model of $Z$ crosstalk considers the closed-system evolution generated by the Hamiltonian
$H=\frac{J}{2}Z_{0}Z_{1}.$ (86)
To monitor how this coherent interaction influences the main qubit we track the time evolution of its Bloch-vector components. Instead of the Schrödinger picture—where the state evolves and operators are static—we adopt the Heisenberg picture, in which operators satisfy
$\frac{d}{dt}O(t)=\mathcal{L}_{H}^{\dagger}[O(t)],\qquad\mathcal{L}_{H}(\cdot)=-\,i[H,\,\cdot\,].$ (87)
Applying $\mathcal{L}_{H}^{\dagger}$ to the Pauli $X$ operator gives
$\mathcal{L}_{H}^{\dagger}(X_{0})$ $=-J\,Y_{0}Z_{1},$ $\mathcal{L}_{H}^{\dagger}(Y_{0}Z_{1})$ $=J\,X_{0},$ (88)
while similarly for Pauli $Y$ operator
$\mathcal{L}_{H}^{\dagger}(Y_{0})$ $=J\,X_{0}Z_{1},$ $\mathcal{L}_{H}^{\dagger}(X_{0}Z_{1})$ $=-J\,Y_{0},$ (89)
We can see that these four operators therefore form a closed set
$\mathcal{O}=\bigl\{X_{0},\;Y_{0}Z_{1},\;Y_{0},\;X_{0}Z_{1}\bigr\},$ (90)
whose dynamics can be written as the linear system
$\dot{\mathbf{v}}(t)=M\,\mathbf{v}(t),\qquad$ (91)
where
$\mathbf{v}=\begin{pmatrix}X_{0}\\ Y_{0}Z_{1}\\ Y_{0}\\ X_{0}Z_{1}\end{pmatrix},\quad\quad M=\begin{pmatrix}0&-J&0&0\\ J&0&0&0\\ 0&0&0&J\\ 0&0&-J&0\end{pmatrix}$. (92)
The formal solution is $\mathbf{v}(t)=e^{Mt}\mathbf{v}(0)$. Because $M$ consists of two identical $2\times 2$ blocks, its eigen-decomposition immediately yields
$\langle X_{0}\rangle(t)$ $=\langle X_{0}\rangle_{0}\cos(Jt)-\langle Y_{0}Z_{1}\rangle_{0}\sin(Jt),$ (93)
$\langle Y_{0}\rangle(t)$ $=\langle Y_{0}\rangle_{0}\cos(Jt)+\langle X_{0}Z_{1}\rangle_{0}\sin(Jt),$ (94)
Consequently, the main-qubit Bloch components $v_{x}(t)=\langle X_{0}\rangle(t)$ and $v_{y}(t)=\langle Y_{0}\rangle(t)$ oscillate at angular frequency $J$, with amplitudes set by the initial two-qubit correlations $\langle Y_{0}Z_{1}\rangle_{0}$ and $\langle X_{0}Z_{1}\rangle_{0}$.
C.2 $Z$ crosstalk between two qubits with dissipator
We now supplement the coherent $Z$ interaction with two Markovian noise channels—amplitude damping (rate $\gamma_{\downarrow q}$, governing the $T_{1}$ time) and pure dephasing (rate $\gamma_{\phi q}$, governing the $T_{2}$ time). With the standard dissipator
$\mathcal{D}[L](\rho)=L\rho L^{\dagger}-\frac{1}{2}\{L^{\dagger}L,\rho\},$ (95)
also we consider the detuning for the main qubit with the Hamiltonian now:
$H=-\frac{\omega_{0}}{2}Z_{0}+\frac{J}{2}Z_{0}Z_{1},$ (96)
the full Lindbladian reads
$\mathcal{L}=\mathcal{L}_{H}+\sum_{q=0}^{1}\gamma_{\downarrow q}\mathcal{D}[\sigma^{-}_{q}]+\gamma_{\phi q}\mathcal{D}[Z_{q}]$ (97)
Action on the operator basis. Introduce the shorthands
$\Gamma_{0}=\tfrac{\gamma_{\downarrow 0}}{2}+2\gamma_{\phi 0},\qquad\Gamma_{1}=\gamma_{\downarrow 1},$
and evaluate $\mathcal{L}^{\dagger}$ on $\mathcal{O}=\{X_{0},Y_{0}Z_{1},Y_{0},X_{0}Z_{1}\}$:
$\mathcal{L}^{\dagger}(X_{0})$ $=-\Gamma_{0}\,X_{0}-J\,Y_{0}Z_{1}+\omega_{0}Y_{0},$
$\mathcal{L}^{\dagger}(Y_{0}Z_{1})$ $=J\,X_{0}-(\Gamma_{0}+\Gamma_{1})\,Y_{0}Z_{1}+\Gamma_{1}\,Y_{0}-\omega_{0}X_{0}Z_{1},$
$\mathcal{L}^{\dagger}(Y_{0})$ $=-\omega_{0}X_{0}-\Gamma_{0}\,Y_{0}+J\,X_{0}Z_{1},$
$\mathcal{L}^{\dagger}(X_{0}Z_{1})$ $=\Gamma_{1}\,X_{0}+\omega_{0}Y_{0}Z_{1}-J\,Y_{0}-(\Gamma_{0}+\Gamma_{1})\,X_{0}Z_{1}.$
Grouping the coefficients gives the linear system $\dot{\mathbf{v}}=M\mathbf{v}$ with
$M=\begin{pmatrix}-\Gamma_{0}&-J&\omega_{0}&0\\ +J&-\Gamma_{0}-\Gamma_{1}&+\Gamma_{1}&-\omega_{0}\\ -\omega_{0}&0&-\Gamma_{0}&+J\\ +\Gamma_{1}&\omega_{0}&-J&-\Gamma_{0}-\Gamma_{1}\end{pmatrix}$, (98)
where we can see that amplitude damping on the spectator qubit ($\Gamma_{1}\neq 0$) couples the two $2\times 2$ off diagonal sub-blocks that were decoupled in the purely coherent case.
Block-diagonalization. Since the contribution of amplitude damping, the matrix $M$ is no longer block-diagonal. We want to find a representation that block diagonalizes $M$ and provides more structure on the problem. Define ladder-operator–like combinations
$\Sigma_{0}^{\pm}=\frac{1}{2}\bigl(X_{0}\mp iY_{0}\bigr),\qquad\Sigma_{1}^{\pm}=\frac{1}{2}\bigl(X_{0}Z_{1}\mp iY_{0}Z_{1}\bigr),$ (99)
and set $\mathbf{w}=C\mathbf{v}$ with
$\mathbf{w}=\begin{pmatrix}\Sigma^{+}_{0}\\ \Sigma^{+}_{1}\\ \Sigma^{-}_{0}\\ \Sigma^{-}_{1}\end{pmatrix},\qquad C=\tfrac{1}{2}\begin{pmatrix}1&0&-i&0\\ 0&-i&0&1\\ 1&0&+i&0\\ 0&+i&0&1\end{pmatrix}$. (100)
In this basis the generator splits into two $2\times 2$ blocks,
$\tilde{M}=C\,M\,C^{-1}=\tilde{M}_{+}\oplus\tilde{M}_{-}$,
$\tilde{M}_{\pm}=\begin{pmatrix}-\Gamma_{0}+i\omega_{0}&\mp\,iJ\\ \mp\,iJ+\Gamma_{1}&-\Gamma_{0}-\Gamma_{1}+i\omega_{0}\end{pmatrix}$.
Note that the $\tilde{M}_{+}$ block completely determines the dynamics of the operator, since $\tilde{M}_{-}$ is simply its complex conjugate. Therefore, it suffices to focus exclusively on the dynamics governed by the $\tilde{M}_{+}$ block.
Exponential of superoperator matrix. To obtain the matrix representation of the propagator $e^{\tilde{M}_{+}t}$, it is convenient to decompose $\tilde{M}_{+}$ into its traceful and traceless parts:
$\tilde{M}_{+}=\bigl(-\Gamma_{0}-\tfrac{\Gamma_{1}}{2}+i\omega_{0}\bigr)I$ $+B,$ (101)
$B=\begin{pmatrix}\tfrac{\Gamma_{1}}{2}&-iJ\\ \Gamma_{1}-iJ&-\tfrac{\Gamma_{1}}{2}\end{pmatrix}$, $\quad\text{Tr}(B)=0$ (102)
so that $\det B=-\Omega^{2}$, where $\Omega=\tfrac{\Gamma_{1}}{2}-iJ$ and $B^{2}=-\Omega^{2}I$. The closed-form expression for $e^{Bt}$ follows from its Taylor series:
$e^{Bt}=\sum_{n=0}^{\infty}\frac{(Bt)^{n}}{n!}=\cosh(\Omega t)\,I+\frac{\sinh(\Omega t)}{\Omega}\,B,$ (103)
and hence
$e^{\tilde{M}_{+}t}=e^{-(\Gamma_{0}+\tfrac{\Gamma_{1}}{2}-i\omega_{0})t}\begin{pmatrix}C+\tfrac{\Gamma_{1}}{2\Omega}\,S&-i\,\tfrac{J}{\Omega}\,S\\[8.5359pt] \tfrac{\Gamma_{1}-iJ}{\Omega}\,S&C-\tfrac{\Gamma_{1}}{2\Omega}\,S\end{pmatrix}$, (104)
where we have introduced the shorthand notation $C=\cosh(\Omega t)$ and $S=\sinh(\Omega t)$.
Initial conditions. Now consider a different initial condition for $\mathbf{w}(0)$. Suppose the main qubit is prepared in the $\lvert$$\rangle$ state (the analyses for $\lvert\pm\rangle$ and $\lvert\pm i\rangle$ are identical), and the spectator qubit is also in the $\lvert$$\rangle$ state. Hence,
$\langle\Sigma^{+}_{0}\rangle_{0}$ $=\tfrac{1}{2},$ $\langle\Sigma^{+}_{1}\rangle_{0}$ $=0\quad\Longrightarrow\quad\mathbf{w}_{+}(0)=\begin{pmatrix}\tfrac{1}{2}\\ 0\end{pmatrix}$. (105)
It follows that
$\langle\Sigma^{+}_{0}\rangle(t)=\tfrac{1}{2}\,e^{-(\Gamma_{0}+\tfrac{\Gamma_{1}}{2}-i\omega_{0})t}\Bigl(\cosh(\Omega t)+\tfrac{\Gamma_{1}}{2\Omega}\,\sinh(\Omega t)\Bigr).$ (106)
The Bloch vector components are then
$v_{x}(t)=2\mathrm{Re}\,\langle\Sigma^{+}_{0}\rangle(t),\quad v_{y}(t)=-2\mathrm{Im}\,\langle\Sigma^{+}_{0}\rangle(t).$
In the limit $\Gamma_{1}/(2J)\ll 1$, we have $\Omega\approx-iJ$, and the Bloch vectors become
$v_{x}(t)$ $\approx e^{-(\Gamma_{0}+\tfrac{\Gamma_{1}}{2})t}\cos(\omega_{0}t)\cos(Jt),$
$v_{y}(t)$ $\approx-e^{-(\Gamma_{0}+\tfrac{\Gamma_{1}}{2})t}\sin(\omega_{0}t)\cos(Jt).$ (107)
We see that the $Z$ crosstalk term induces an additional oscillation at frequency $J$. In the opposite limit $\Gamma_{1}/(2J)\gg 1$, where $\Omega\approx\tfrac{\Gamma_{2}}{2}$, one finds
$v_{x}(t)$ $\approx e^{-(\Gamma_{0}+\Gamma_{1})t}\cos(\omega_{0}t),$
$v_{y}(t)$ $\approx-e^{-(\Gamma_{0}+\Gamma_{1})t}\sin(\omega_{0}t).$ (108)
In this regime, the $Z$ crosstalk contribution is negligible, and only purely damped oscillations are observed.
For the initial condition with main qubit in $|+\rangle$ state and spectator in the $|0\rangle$ state, we have:
$\langle\Sigma^{+}_{0}\rangle_{0}$ $=\tfrac{1}{2},$ $\langle\Sigma^{+}_{1}\rangle_{0}$ $=\tfrac{1}{2}\quad\Longrightarrow\quad\mathbf{w}_{+}(0)=\begin{pmatrix}\tfrac{1}{2}\\ \tfrac{1}{2}\end{pmatrix}$. (109)
Since
$\langle\Sigma_{0}^{+}\rangle(t)$ $=\tfrac{1}{2}\,e^{-(\Gamma_{0}+\tfrac{\Gamma_{1}}{2})t}e^{i\omega_{0}t}\Bigl[\bigl(C+\tfrac{\Gamma_{1}}{2\Omega}S\bigr)+\bigl(-\,i\tfrac{J}{\Omega}S\bigr)\Bigr]$
$=\tfrac{1}{2}\,e^{-(\Gamma_{0}+\tfrac{\Gamma_{1}}{2})t}e^{i\omega_{0}t}\bigl(C+S)$
$=\tfrac{1}{2}\,e^{-\Gamma_{0}t}\,e^{\,i(\omega_{0}-J)t},$ (110)
where we used $C+S=e^{\Omega t}$ and $\Omega=\tfrac{\Gamma_{1}}{2}-iJ$. Hence the Bloch components
$v_{x}(t)$ $=2\text{Re}\langle\Sigma_{0}^{+}\rangle(t)=e^{-\Gamma_{0}t}\cos((\omega_{0}-J)t),$
$v_{y}(t)$ $=-2\text{Im}\langle\Sigma_{0}^{+}\rangle(t)=e^{-\Gamma_{0}t}\sin((\omega_{0}-J)t).$ (111)
where the amplitude damping effect ($\Gamma_{1}$) is completely disappear of since the spectator qubit is already in the ground state $|0\rangle$.
C.3 $Z$ crosstalk with NN spectator qubits
We now generalize the single-spectator case to NN spectators. Consider the $Z$ crosstalk Hamiltonian:
$H=-\frac{\omega_{0}}{2}Z_{0}+\sum_{q=1}^{N}\frac{J_{0q}}{2}Z_{0}Z_{q},\quad\mathcal{L}^{\dagger}_{H}(\cdot)=i[H,\,\cdot\,],$ (112)
and the dissipators
$\sum_{q=0}^{N}\gamma_{\downarrow q}\mathcal{D}[\sigma_{q}^{-}]+\gamma_{\phi q}\mathcal{D}[Z_{q}]$ (113)
Close operator set. A key observation is that the operators
$\mathcal{O}$ $=\{X_{0},Y_{0}\}\otimes\{I,Z_{1}\}\otimes\cdots\otimes\{I,Z_{N}\}$
$=\bigl\{X_{0}Z_{1}^{s_{1}}Z_{2}^{s_{2}}\cdots Z_{N}^{s_{N}},\;Y_{0}Z_{1}^{s_{1}}Z_{2}^{s_{2}}\cdots Z_{N}^{s_{N}}\bigr\},$ (114)
with $\mathbf{s}=(s_{1},\dots,s_{N})\in\{0,1\}^{N}$, form a closed set under the Lindbladian. We then define the ladder operators
$\Sigma_{\mathbf{s}}^{\pm}=\tfrac{1}{2}\,(X_{0}\mp iY_{0})\,Z_{1}^{s_{1}}Z_{2}^{s_{2}}\cdots Z_{N}^{s_{N}}.$ (115)
Let us examine how each superoperator acts on the elements of this set. For the qubit detuning and $Z$ crosstalk terms:
$i[Z_{0},\Sigma_{\mathbf{s}}^{\pm}]$ $=\mp 2i\Sigma_{\mathbf{s}}^{\pm},$
$i\bigl[Z_{0}Z_{q},\Sigma_{\mathbf{s}}^{\pm}\bigr]$ $=\mp 2i\,\Sigma_{\mathbf{s}\oplus\mathbf{e}_{q}}^{\pm}.$ (116)
Note that the factor of $\mp 2$ arises from the commutation relation $Z,\sigma^{\pm}]=\mp 2\,\sigma^{\pm}$, and the modular addition follows from $Z_{q}Z_{q}^{s_{q}}=Z_{q}^{s_{q}\oplus 1}$. For amplitude damping:
$\mathcal{D}^{\dagger}[\sigma_{q}^{-}]\bigl(\Sigma_{\mathbf{s}}^{\pm}\bigr)=\begin{cases}-\tfrac{1}{2}\,\Sigma_{\mathbf{s}}^{\pm},&q=0,\\[5.69046pt] \delta_{s_{q},1}\bigl(\Sigma_{\mathbf{s}\oplus\mathbf{e}_{q}}^{\pm}-\Sigma_{\mathbf{s}}^{\pm}\bigr),&q\geq 1.\end{cases}$ (117)
For pure dephasing:
$\mathcal{D}^{\dagger}[Z_{q}]\bigl(\Sigma_{\mathbf{s}}^{\pm}\bigr)=\begin{cases}-2\,\Sigma_{\mathbf{s}}^{\pm},&q=0,\\[5.69046pt] 0,&q\geq 1.\end{cases}$ (118)
We immediately see that the Lindbladian maps each element of $\mathcal{O}_{+}=\{\Sigma_{\mathbf{s}}^{+}\}_{\mathbf{s}\in\{0,1\}^{N}}$ and $\mathcal{O}_{-}=\{\Sigma_{\mathbf{s}}^{-}\}_{\mathbf{s}\in\{0,1\}^{N}}$ to itself. Therefore, both $\mathcal{O}_{+}$ and $\mathcal{O}_{-}$ form closed operator sets under $\mathcal{L}$. Moreover, this implies that the matrix $\tilde{M}$ is block-diagonal in this basis.
Encoding of the NN-qubit parity Hilbert space. To better inspect the structure of the superoperator matrix $\tilde{M}$, we define the computational basis
$|\mathbf{s}\rangle$ $=|s_{1}\rangle\otimes\cdots\otimes|s_{N}\rangle,$ $s_{k}$ $\in\{0,1\}.$ (119)
For each closed operator set $\mathcal{O}_{\pm}$, we then identify
$\Sigma^{+}_{\mathbf{s}}$ $\longleftrightarrow|\mathbf{s}\rangle\in\mathscr{H}_{+},$ $\Sigma^{-}_{\mathbf{s}}$ $\longleftrightarrow|\mathbf{s}\rangle\in\mathscr{H}_{-}.$ (120)
Thus, $\mathscr{H}_{+}\cong(\mathbb{C}^{2})^{\otimes N}\cong\mathscr{H}_{-}$. On this parity Hilbert space, we define the corresponding Pauli operators by
$Z_{k}\,|\mathbf{s}\rangle$ $=(-1)^{s_{k}}\,|\mathbf{s}\rangle,$ $X_{k}\,|\mathbf{s}\rangle$ $=|\mathbf{s}\oplus\mathbf{e}_{k}\rangle.$ (121)
Note that these Pauli operators act on the abstract parity Hilbert space and differ from the physical Pauli operators introduced earlier.
Now we describe the form of the superoperator in this representation. First, consider the ZZ‐crosstalk term:
$\mathcal{L}_{H}^{\dagger}\mathrel{\mathop{\ordinarycolon}}\Sigma_{\mathbf{s}}^{\pm}\mapsto\pm i\omega_{0}\Sigma_{\mathbf{s}}^{\pm}\mp\sum_{q=1}^{N}iJ_{0q}\,\Sigma_{\mathbf{s}\oplus\mathbf{e}_{q}}^{\pm},$
$\implies\tilde{M}_{\pm}^{(H)}=\pm i\omega_{0}I\mp\sum_{q=1}^{N}iJ_{0q}\,X_{q}.$ (122)
Next, for the amplitude‐damping dissipator:
$\sum_{q=0}^{N}\gamma_{\downarrow q}\,\mathcal{D}^{\dagger}[\sigma_{q}^{-}]$ $\mathrel{\mathop{\ordinarycolon}}\Sigma_{\mathbf{s}}^{\pm}\mapsto-\frac{\gamma_{\downarrow 0}}{2}\,\Sigma_{\mathbf{s}}^{\pm}+\sum_{q=1}^{N}\gamma_{\downarrow q}\delta_{s_{q},1}\Bigl(\Sigma^{\pm}_{\mathbf{s}\oplus\mathbf{e}_{q}}-\Sigma_{\mathbf{s}}^{\pm}\Bigr),$
$\implies\tilde{M}^{(\downarrow)}_{\pm}$ $=-\frac{\gamma_{\downarrow 0}}{2}I+\sum_{q=1}^{N}\gamma_{\downarrow q}\,\Big(\sigma_{q}^{-}-\frac{1}{2}(I-Z_{q})\Big).$ (123)
and the dephasing dissipator:
$\sum_{q=0}^{N}\gamma_{\phi q}\mathcal{D}^{\dagger}[Z_{q}]\mathrel{\mathop{\ordinarycolon}}\Sigma_{\mathbf{s}}^{\pm}\mapsto-2\gamma_{\phi 0}\Sigma_{\mathbf{s}}^{\pm},$
$\implies\tilde{M}_{\pm}^{(\phi)}=-2\gamma_{\phi 0}I.$ (124)
Each of the above superoperator matrices is a sum of single‐qubit operators; hence the overall superoperator matrix $\tilde{M}_{\pm}$ can be written as a Kronecker sum:
$\tilde{M}_{\pm}$ $=\tilde{M}^{(H)}_{\pm}+\tilde{M}^{(\downarrow)}_{\pm}+\tilde{M}^{(\phi)}_{\pm}$
$=-\Big(2\gamma_{\phi 0}+\sum_{q=0}^{N}\frac{\gamma_{\downarrow q}}{2}\pm i\omega_{0}\Big)I$
$\hskip 28.45274pt+\sum_{q=1}^{N}I^{\otimes(q-1)}\otimes B^{(q)}_{\pm}\otimes I^{\otimes(N-q)},$ (125)
where
$B^{(q)}_{\pm}$ $=\mp iJ_{0q}X_{q}+\gamma_{\downarrow q}\Big(\sigma_{q}^{-}+\frac{1}{2}Z_{q}\Big)$
$=\begin{pmatrix}\tfrac{\Gamma_{q}}{2}&\mp iJ_{0q}\\ \mp iJ_{0q}+\Gamma_{q}&-\tfrac{\Gamma_{q}}{2}\end{pmatrix}$, (126)
with $\Gamma_{0}=2\gamma_{\phi 0}+\tfrac{\gamma_{\downarrow 0}}{2}$ and $\Gamma_{q}=\gamma_{\downarrow q}$ for $q\geq 1$. We immediately see that this matrix is identical to the one‐spectator case. Moreover, the propagator can be written as
$e^{\tilde{M}_{\pm}t}=e^{(-\Gamma_{0}+i\omega_{0})t}\,\bigotimes_{q=1}^{N}\bigl(e^{-\tfrac{\Gamma_{q}t}{2}}\,e^{B^{(q)}_{\pm}t}\bigr).$ (127)
Note that although the number of element of the closed operator set scales exponentially with the number of spectator qubits, the propagator admits an elegant tensor‐product decomposition into single‐qubit matrices.
Initial conditions. We now consider two different initial conditions for the main and spectator qubits.
Suppose the main qubit is prepared in the $\lvert$$\rangle$ state and each spectator qubit is likewise in the $\lvert$$\rangle$ state, i.e. $\lvert+\rangle^{\otimes N}$. Hence,
$\langle\Sigma_{0}^{+}\rangle_{0}$ $=\tfrac{1}{2},$ $\langle\Sigma_{\mathbf{s}}^{+}\rangle_{0}$ $=0\quad\text{for }\mathbf{s}\neq 0,$ $\Longrightarrow\quad\mathbf{w}_{+}(0)$ $=\begin{pmatrix}\tfrac{1}{2}\\ 0\\ \vdots\\ 0\end{pmatrix}$.
It is often more convenient to use Dirac notation:
$\mathbf{w}_{+}(0)=\tfrac{1}{2}\,\lvert 0\rangle^{\otimes N}.$
The first component of $\mathbf{w}_{+}(t)$ is then
$\langle\Sigma_{0}^{+}\rangle(t)$ $=\bigl[e^{\tilde{M}_{+}t}\,\mathbf{w}_{+}(0)\bigr]_{0}$
$=\tfrac{1}{2}e^{(-\Gamma_{0}+i\omega_{0})t}\Big[\bigotimes_{q=1}^{N}e^{-\tfrac{\Gamma_{q}t}{2}}e^{B^{(q)}t}|0\rangle^{\otimes N}\Big]_{0}$
$=\tfrac{1}{2}\,e^{(-\Gamma_{0}+i\omega_{0})\,t}\prod_{q=1}^{N}e^{-\tfrac{\Gamma_{q}\,t}{2}}\langle 0\rvert e^{B_{+}^{(q)}t}\lvert 0\rangle.$ (128)
The matrix element is
$\langle 0\rvert e^{B_{+}^{(q)}t}\lvert 0\rangle=C_{q}+\frac{\Gamma_{q}}{2\,\Omega_{q}}\,S_{q},$
where $C_{q}=\cosh(\Omega_{q}t)$, $S_{q}=\sinh(\Omega_{q}t)$, and $\Omega_{q}=\tfrac{\Gamma_{q}}{2}-iJ_{0q}$.
In the limit $\Gamma_{q}/(2J_{0q})\ll 1$, one has $\langle 0\rvert e^{B_{+}^{(q)}t}\lvert 0\rangle\approx\cos(J_{0q}t)$, and therefore
$v_{x}(t)$ $\approx e^{-\Gamma_{0}t}\cos(\omega_{0}t)\prod_{q=1}^{N}e^{-\tfrac{\Gamma_{q}t}{2}}\cos(J_{0q}t),$ (129)
$v_{y}(t)$ $\approx-e^{-\Gamma_{0}t}\sin(\omega_{0}t)\prod_{q=1}^{N}e^{-\tfrac{\Gamma_{q}t}{2}}\cos(J_{0q}t).$ (130)
In the opposite limit $\Gamma_{q}/(2J_{0q})\gg 1$, one has $\langle 0\rvert e^{B_{+}^{(q)}t}\lvert 0\rangle\approx e^{-\tfrac{\Gamma_{q}t}{2}}$, giving
$v_{x}(t)$ $\approx e^{-\sum_{q=0}^{N}\Gamma_{q}\,t}\cos(\omega_{0}t),$ (131)
$v_{y}(t)$ $\approx-e^{-\sum_{q=0}^{N}\Gamma_{q}\,t}\sin(\omega_{0}t).$ (132)
Now suppose each spectator qubit is in the $\lvert 0\rangle$ state, i.e. $\lvert 0\rangle^{\otimes N}$. Hence,
$\langle\Sigma_{0}^{+}\rangle_{0}=\tfrac{1}{2},\quad\langle\Sigma_{\mathbf{s}}^{+}\rangle_{0}=\tfrac{1}{2}\quad\forall\,\mathbf{s},\quad\Longrightarrow\quad\mathbf{w}_{+}(0)=\begin{pmatrix}\tfrac{1}{2}\\ \tfrac{1}{2}\\ \vdots\\ \tfrac{1}{2}\end{pmatrix}$.
In this case,
$\mathbf{w}_{+}(0)=2^{\tfrac{N}{2}-1}\,\lvert+\rangle^{\otimes N},$
and the corresponding amplitude is
[…]
As in the single‐spectator case,
$\sqrt{2}\,\langle 0\rvert e^{B^{(q)}t}\lvert+\rangle=e^{-\tfrac{\Gamma_{q}\,t}{2}}\,e^{-iJ_{0q}t}.$
Therefore,
$v_{x}(t)$ $=e^{-\sum_{q=0}^{N}\Gamma_{q}\,t}\,\cos\bigl(\omega_{0}t-\sum_{q=1}^{N}J_{0q}t\bigr),$ (133)
$v_{y}(t)$ $=-\,e^{-\sum_{q=0}^{N}\Gamma_{q}\,t}\,\sin\bigl(\omega_{0}t-\sum_{q=1}^{N}J_{0q}t\bigr).$ (134)
General pure product initial states. We generalize the result to arbitrary pure product states for the main and spectator qubits. Consider the single‐qubit states
$|\psi_{q}\rangle$ $=\cos\tfrac{\theta_{q}}{2}\,|0\rangle+e^{i\phi_{q}}\sin\tfrac{\theta_{q}}{2}\,|1\rangle,$ (135)
so that the global initial state is
$|\Psi(0)\rangle=\bigotimes_{q=0}^{N}|\psi_{q}\rangle.$ (136)
Because the global state factorizes, the initial expectation values also factorize:
$\langle\Sigma_{\mathbf{s}}^{+}\rangle_{0}$ $=\tfrac{1}{2}\,\langle\psi_{0}|(X_{0}-iY_{0})|\psi_{0}\rangle\prod_{q=1}^{N}\langle\psi_{q}|Z_{q}^{s_{q}}|\psi_{q}\rangle.$ (137)
For the main qubit,
$\tfrac{1}{2}\,\langle\psi_{0}|(X_{0}-iY_{0})|\psi_{0}\rangle$ $=\tfrac{1}{2}\,(x_{0}-iy_{0}),$ (138)
and for each spectator qubit,
$\langle\psi_{q}|Z_{q}^{s_{q}}|\psi_{q}\rangle$ $=\left\{\begin{array}[]{ll}1,&s_{q}=0,\\ z_{q},&s_{q}=1.\end{array}\right.=z_{q}^{s_{q}},$ (141)
Hence, the component labeled by $\mathbf{s}$ is
$\langle\Sigma_{\mathbf{s}}^{+}\rangle_{0}=\tfrac{1}{2}\,(x_{0}-iy_{0})\,\prod_{q=1}^{N}z_{q}^{s_{q}}.$ (142)
It follows that
$\mathbf{w}_{+}(0)$ $=\tfrac{1}{2}\,(x_{0}-iy_{0})\sum_{\mathbf{s}\in\{0,1\}^{N}}\Bigl(\prod_{q=1}^{N}z_{q}^{s_{q}}\Bigr)|\mathbf{s}\rangle$
$=\tfrac{x_{0}-iy_{0}}{2}\bigotimes_{q=1}^{N}\bigl(|0\rangle+z_{q}\,|1\rangle\bigr).$ (143)
Thus, in the parity Hilbert space,
$\mathbf{w}_{+}(0)=\tfrac{x_{0}-iy_{0}}{2}\bigotimes_{q=1}^{N}|w_{q}\rangle,\quad|w_{q}\rangle=|0\rangle+z_{q}\,|1\rangle.$ (144)
Finally, the main‐qubit Bloch‐vector dynamics follow from
$\langle\Sigma_{0}^{+}\rangle(t)$ $=\bigl[e^{\tilde{M}_{+}t}\,\mathbf{w}_{+}(0)\bigr]_{0}$
$=e^{(-\Gamma_{0}+i\omega_{0})\,t}\tfrac{x_{0}-iy_{0}}{2}\prod_{q=1}^{N}e^{-\tfrac{\Gamma_{q}\,t}{2}}\langle 0|e^{B_{+}^{(q)}t}|w_{q}\rangle$
$=e^{(-\Gamma_{0}+i\omega_{0})\,t}\tfrac{x_{0}-iy_{0}}{2}\prod_{q=1}^{N}e^{-\tfrac{\Gamma_{q}\,t}{2}}\left(C_{q}+\tfrac{\tfrac{\Gamma_{q}}{2}-iz_{q}J}{\Omega_{q}}S_{q}\right)$ (145)